A solution containing 121.8 g of Zn(NO3)2 per litre has a density of 1.107 g/mL.

Calculate its molal concentration.


1
Expert's answer
2021-08-16T02:45:34-0400

Molal concentration of solution is 0.875 mol/kg


Cm =moles of solute/kg of solvent


V(solute)=121.8 g/1.107g/mL-1=116.35mL


V(solvent)= 1000mL-116.35mL=883.65 mL=0.88L=0.88 kg


n=m/M=121.8g/157.4g/mol=0.77 mol


Cm=0.77 mol/0.88 kg=0.875 mol/kg